2. x = cos 20 y = sin 40 a. r = 11

How do I find the equation of the tangent line?
Answer #1

This is the proper solution of your question.

Thank you.

2 x = (os (20) dx = -28in (20) y = din (40) dy = 4 (os (40) do do lope dy = - 4 Cos (40) = -2 Cos (40) dx & Sin (20) Sin (20)

(a) r= 1 = 0 dy y o = dx -2 Cos (II) = -2(-1) = 2 4-0 = 2(x-o) Ty=2x (5) r = t = 0 dy = -2 Cos (20) da Sin (II) CTT